OpenStudy (anonymous):

Find all solutions to the equation. cos2x + 2 cos x + 1 = 0

OpenStudy (anonymous):

rewrite \(\cos(2x)=2\cos^2(x)-1\) and solve that one

OpenStudy (anonymous):

in other words, solve \[2\cos^2(x)-1+\cos(x)+1=0\]

OpenStudy (anonymous):

can you explain how to solve it?

OpenStudy (anonymous):

first step would be to note that \(1-1=0\) and solve \[2\cos^2(x)+\cos(x)=0\] then factor as \(\cos(x)\left(2\cos(x)+1\right)=0\) finally solve \[\cos(x)=0\] and also \[2\cos(x)+1=0\] \[2\cos(x)=-1\] \[\cos(x)=-\frac{1}{2}\]

OpenStudy (anonymous):

thank you so much! :)